OpenStudy (anonymous):

A ball is thrown into the air. A sensor measures the height of the ball above the ground. the height of the ball is described by the equation: y=-5x(x-4), where x represents the time in seconds. determine the intervals of time when the ball is above a height of 15meters.

OpenStudy (anonymous):

will fan and medal for answer

OpenStudy (anonymous):

There are two options with this equation, one is you can graph it on a graphing calculator and find what x is when you put 15 for y, but since most graphing calculators don't let you input for a zero for y solve it like this. Set 15=y because it asks for what the time (x) is when the ball is 15 m above the ground. So your new equation is 15= -5x(x-4) now simplify 15=-5x^2+20x subtract 15 from both sides -5x^2+20x-15=0 factor the equation (5x+15)(x-1)=0 Then solve for each factored part for x 5x+15=0 5x=-15 x=-3 x-1=0 x=1 since -3 seconds is not possible in our world, x= 1 so after 1 second the ball will reach 15 meters. If you put in y=-5x^2+20 into a graphing calculator you should get the same answer. Good Luck, Hope I helped :)

OpenStudy (anonymous):

Could you fan? @Twiztidsplifff

OpenStudy (anonymous):

nitely

OpenStudy (anonymous):

I think @I_Need_Help_With_Home is correct, she explained it really good